A 0.500 L solution contains 1.68 g of NH3(aq) and 4.05 g of NH4CI(aq). Kb=1.80*10^-5

a.determine the pH of this solution.

b.determine the volume of 6.0 M HCI(aq) that must be added to the above buffer solution to change the pH to9.00
Answer #1

a)pOH = pKb + log [BH+]/[B],

pH = 14.0 - pOH

[BH+] = moles of nh4cl / vol in litre

moles of nh4cl = 4.05/53.49 = 0.075

[BH+] = 0.151 M

[B] = (1.68/17)/0.500 = 0.197 M

pKb = - log[kb]

pOH = 4.74 + (-0.11) = 4.63

pH = 14-4.63 = 9.37